There is a photograph in an ad and you have to increase the length and width of the photograph by the same amount in order to double the area of the photograph. If the photongraph was 12cm by 8cm. What are the enlarged demensions of the photograph. Using factoring i
Answer by jorel1380(3719)

Let w be the area to be expanded to. Then:
(12+2w)(8+2w)=96*2=192
96+40w+4w²=192
w²+10w-24=0
(w+12)(w-2)=0
w=2 or -12
Using the positive value of 2 as the border, our new dimensions are 16cm.x 12cm. ☺☺☺☺